# Blind Spot Navigation in LLM Reasoning with Thought Space Explorer

### v2

@@ 3.4 Collaborative Reasoning
3.4.1 Collaborative Weighted Summation
This strategy applies for gradient-available LLMs. We encompass all thought chains in refined thought structure $\mathcal{S}^{\prime}$. First, we use gradient information to recalculate and select key nodes of each chain. For each key node $T_{i_{key}}$, we calculate a weight $w_{i_{key}}$ based on its impact on the solution:
@@ 3.4.1 Collaborative Weighted Summation
where $L_{i_{key}}$ represents the self-information loss at node $T_{i_{key}}$, which reflects the model’s confidence and potential errors in its prediction at that node. The contribution of each key node, denoted as $v_{i_{key}}$, is the value vector that quantifies the influence of the node on the overall reasoning process. This vector can include various factors such as the relevance of the node’s content to the task or the accuracy of its inference. Then for reasoning task $\mathcal{Q}$, we compute the collaborative reasoning score by summing the contributions of all key nodes across all chains:
$C(\mathcal{Q})=\sum_{i=1}^{N}\sum_{j=1}^{K_{i}}w_{i_{key}}\cdot v_{ij}$ (18)
This process integrates the weighted contributions for a reasoning score that involves the individual node’s direct impact and its significance within the context of the entire chain. The final decision $D$ for the task $\mathcal{Q}$ is selected based on the highest collaborative reasoning score:
@@
3.4.2 LM-as-a-Judge
This strategy is suitable for reasoning tasks that require detailed interpretation and judgment, applicable to both gradient-available models and black-box models. In this strategy, the model $\mathcal{L}$ acts as a judge to score each thought chain’s output based on its assessment of reasoning coherence, prediction confidence, or relevance to the task. Based on the evaluation, the model employs a voting mechanism to determine the final output:
$\text{Score}_{i}=\text{Evaluate}_{\mathcal{L}}(C_{q})$ (20)
$D=\arg\max_{q\in\mathcal{Q}}C(q)$ (21)
@@ 3.4.2 LM-as-a-Judge
4 Experiments
We conduct experiments on three reasoning tasks that require mathematical, hierarchical, and comprehensive reasoning abilities to evaluate the effectiveness of our TSE method.
4.1 Experimental Settings
Supplemental Details. We conduct the experiments utilizing the GPT-4o-mini OpenAI (2024) and Llama-3.1-8B-Instruct Hugging Face (2024). Unless otherwise specified, we generate five parallel thought chains with a depth of 5 for each question and use this structure as the basis for new node generation. The temperature for all models is set to the default value of 0.7, with a maximum token limit of 50. All tasks are performed on an NVIDIA 4090 GPU.
Baselines. We apply our method to the simplest multi-chain thought structure and compare the results with several baseline methods, including CoT, Vanilla CoT-SC, ToT, and RATT. This comparison aims to illustrate how our method enhances the thought structure compared to existing approaches. Detailed information is provided in Appendix A.1.
Task Game of 24 Mini Crossword Creative Writing
Success Rate (%) Success by Letter (%) Success by Game (%) Soundness Innovation Coherence Expression Overall
CoT 13.3 38.4 1.8 5.26 5.24 5.03 4.95 5.12
CoT-SC 46.7 52.2 7.3 5.41 5.17 5.17 5.16 5.23
ToT 52.7 75.5 18.7 5.90 5.36 5.40 5.43 5.52
RATT 41.3 79.2 25.5 6.02 6.22 5.69 5.67 5.90
TSE 74.0 82.4 24.1 6.19 6.55 5.74 5.64 6.03
@@ 4.1 Experimental Settings
Table 1: Overall performance on three reasoning tasks based on GPT-4o-mini. Bold and underline indicate the best and the second-best results.
@@
Task Description. We evaluate TSE and the baseline methods on three reasoning datasets with specific tasks: (1) Game of 24, a mathematical challenge whose objective is to use the four basic arithmetic operations to make four given numbers equal 24. The task requires the language models to combine multiple operations to achieve a target outcome, which can evaluate arithmetic reasoning and logical problem-solving capabilities. (2) Mini Crosswords, a game of 5×5 mini crosswords and each input includes the 5 horizontal and 5 vertical clues. To solve this task, the model requires deeper exploration and strategic integration of linguistic clues, allowing us to understand how effectively the evaluated model can expand traditional solution paths and uncover new insights within a complex search space; (3) Creative Writing, a task to construct a coherent passage with four paragraphs, each ending with one of four given sentences.This task compels LLMs to generate imaginative text withlogically sound, and contextually rich.
4.2 Overall Performance
@@
The overall performance is reported in Table 1. We can see our TSE consistently outperforms compared baseline methods across different tasks on most of the metrics. The effectiveness of TSE is attributed to expanding thought structures and exploring solution space to contribute to different aspects of reasoning. Then, we analyze each task separately as follows.
@@
Game of 24. The evaluation of Game of 24 is shown in Table 1, where the TSE-refined CoT-SC method significantly outperforms other thought structures, including vanilla CoT-SC, with an improvement of $58.56\%$, ToT by $40.50\%$, and RATT by $79.04\%$. Additionally, our enhancement of the basic CoT-SC structure achieves accuracy on GPT-4o-mini that matches the performance of the more complex ToT ($b=5$) implemented on GPT-4 Yao et al. (2024). These results highlight the substantial improvements in reasoning capabilities brought about by our method.
@@
Mini Crossword. In our task setup, we evaluate performance based on two metrics: the proportion of correct letters (out of 25 per game) and the proportion of successfully solved games. As shown in Table 1, the CoT-SC with TSE achieves an impressive accuracy rate of $82.4\%$, significantly outperforming vanilla CoT-SC (by $30.2\%$), surpassing ToT by $6.9\%$. However, for the Game metric, RATT performs slightly better than our method, as it can leverage external knowledge through RAG. Specifically, vanilla CoT-SC suffers from limited exploration of thought paths, which reduces its ability to consistently generate accurate answers. By refining CoT-SC with our method, we enhance its depth and diversity in exploring possible solutions, leading to more accurate and coherent outputs. These results further demonstrate the substantial improvements our method brings to enhance language models’ reasoning capabilities.
@@
Creative Writing. As shown in Table 1, the evaluation criteria included Soundness, Innovation, Content Coherence of Reasoning, and Clarity of Expression, applying consistently across all evaluations. We rate each metric on a scale from 0 to 10 in 0.5-point increments, and calculate the overall performance score as the average of these four dimensions. For detailed figure, refer to Appendix A.4.
4.3 Ablation Study
4.3.1 Key Node Selection
@@
To validate the importance of key node selection, we compare (1) the original CoT-SC, (2) the random selection of one node per chain for generation, and (3) our key node selection method. As shown in Table 3, our method significantly improves on Game of 24 compared to original and random approaches. In Creative Writing, particularly for innovation, our method also demonstrates notable enhancements, proving its effectiveness in selecting nodes for generation. The superior performance of our key node selection method is primarily due to its ability to identify nodes containing crucial information for problem-solving. The model only utilizes the most influential information and thus avoids redundancy. In this way, our method enhances the model’s efficiency and creativity.
Game of 24 Creative Writing
Method Success Rate (%) Overall Coherence
@@
Random 9.52 4.82 4.85
Layer-based 6.12 4.84 4.88
TSE 16.32 4.91 4.94
@@
Majority-Vote 46.7 5.81 5.87
Random Sampling 27.3 5.90 5.89
New Chains Only 64.0 5.95 5.87
TSE 74.0 6.03 5.90
@@
4.3.2 Connection Node Selection
We then test the necessity of connection node selection by comparing (1) randomly selecting key nodes to connect, (2) selecting the key nodes based on lower inference layers, and (3) our method. As shown in Table 3, our method significantly superior random selection and layer-based selection on overall reasoning performance, particularly in coherence. This is because random selection lacks specificity and cannot guarantee that the chosen nodes will effectively support the content of new nodes. Although layer-based selection considers the structural hierarchy of information, it does not necessarily reflect the actual importance or applicability of the information. Our method analyzes relationships and reasoning paths between nodes, thus maintaining and strengthening the information flow’s coherence and depth.
4.3.3 Collaborative Reasoning
In this experiment, we continue to test the importance of collaboration method in TSE. We compare four collaboration methods: (1) aggregating answers through majority voting, (2) randomly sampling partial thought chains, (3) using only the outcomes from newly generated chains, and (4) implementing the TSE collaborative method. As shown in Table 3, TSE consistently outperforms the other methods in both tasks. Meanwhile, using only new chains also achieves a high success rate in Game of 24, demonstrating their potential to offer innovative solutions. However, since these new chains lack integration with original chain information, their performance is still weaker than that of TSE’s comprehensive collaboration methods. In Creative Writing, using only new chains proved better than majority voting. This indicates that while new chains can provide innovative content, they may fall short in coherence and depth without sufficient integration with the original chains.
4.4 Strategy Analysis
4.4.1 Impact of Key Node Selection Methods
In this experiment, we compare the impact of our two key node selection methods on model reasoning performance. Here we contrast the gradient-based and self-prompting selection on Llama’s CoT-SC structures. As shown in Figure 3, both methods perform similarly in task Game of 24, while in Creative Writing, the self-prompting method shows superior results. This suggests that both methods are capable of selecting appropriate key nodes. Meanwhile, the self-prompting methods guide the model to review known information as well as consider how to connect different thought fragments. As both methods select the key node of the chain, self-promoting explores blind spots that the model might overlook in the usual autoregressive generation process. By delving deeper into these spaces, the model generates more reliable information from the selected nodes and provides more creative content when generating new nodes. Thus, for tasks requiring high creativity and diverse thinking, a potential strategy is to prompt the model to break free from conventional patterns and generate innovative insights.
Figure 3: Performance of key node selection methods.
4.4.2 Comparison of Connection Node Selection Methods
In this experiment, we compare the impact of two different methods for selecting connection nodes for newly generated nodes on reasoning performance. We evaluate the gradient comparison and semantic relevance methods on Llama-3.1-8B. As shown in Table 3, selecting nodes based on semantic relevance leads to better performance on Game of 24 reasoning and content coherence of reasoning in Creative Writing. The difference possibly arises from two main reasons: (1) Nodes with larger relative gradients do not necessarily have a stronger influence on the content when generating new nodes. (2) Compared to gradient-based selection, the semantic relevance method more effectively captures the actual content and meaning relationships between nodes, leading to more suitable connection points for new nodes. This demonstrates that considering semantic relevance when selecting connection points provides more effective support for reasoning performance.
@@ 4.4.2 Comparison of Connection Node Selection Methods
4.4.3 Comparison of Collaborative Reasoning Methods
In this experiment, we compare the performance of CWS and LM-as-a-Judge. As shown in Table 3, in task Game of 24, which requires dense logic and has well-defined objectives, the LM-as-a-Judge method accurately and quickly identifies the correct answers by directly evaluating each thought chain and selecting the optimal solution. This method relies on the model’s ability to assess individual outputs, allowing for rapid selection of the best solution—especially useful in tasks that require precise calculations and quick responses. On the other hand, in task Creative Writing, which requires innovation and diversity, CWS evaluates and weights the contributions of each thought chain. The final output represents a blend of multiple perspectives and enhances the richness and depth of the narrative significantly. We further analyze the thought chain efficacy and provide a case study to illustrate the effectiveness of TSE generation and collaboration, referring to Appendix A.3.
5 Conclusion
@@
Limitations
We develop a framework with multiple strategies to enhance thought structures for LLMs, where the expansion of thought structures through self-generated nodes might lead to over-fitting to existing patterns as the process works without integrating external knowledge. This may limit the TSE method’s ability to explore more diverse thought patterns. Besides, the experiments are conducted on only two language models, which may not provide a comprehensive view of the TSE’s generalization capability across LLMs with varying sizes and pre-training processes. Moreover, the TSE is mainly evaluated on specific tasks and these tasks cannot fully reflect the complexities of real-world scenarios where reasoning tasks can be variable and with more complex and external solution spaces. We left these potential explorations as future work.
@@ A.1 Baselines
A.2 Task Description
We evaluate TSE and the baseline methods on three reasoning datasets with specific tasks: (1) Game of 24, a mathematical challenge whose objective is to use the four basic arithmetic operations to make four given numbers equal 24. The task requires the language models to combine multiple operations to achieve a target outcome, which can evaluate arithmetic reasoning and logical problem-solving capabilities. (2) Mini Crosswords, a game of 5×5 mini crosswords and each input includes the 5 horizontal and 5 vertical clues. The expected output is a completed $5×5$ crossword board containing 25 letters. To solve this task, the model requires deeper exploration and strategic integration of linguistic clues, allowing us to understand how effectively the evaluated model can expand traditional solution paths and uncover new insights within a complex search space; (3) Creative Writing, a task to construct a coherent passage with four paragraphs, each ending with one of four given sentences.This task compels LLMs to generate imaginative text withlogically sound, and contextually rich. To ensure a rigorous and comprehensive assessment, we conduct half of the evaluations using GPT-4 OpenAI (2023), while the other half are performed by a panel of expert human annotators. In this task, we select 100 open-ended questions from several prompts listed on Reedsy. com [^2] https://blog.reedsy.com/creative-writing-prompts/#::̃text=When%20the%20idea%20to%20start%20a%20weekly%20newsletter as our input.
A.3 Thought Chain Efficacy
@@
A.4 Creative Writing

### v3

@@ Introduction
Collaborative Weighted Summation
This strategy applies to gradient-available LLMs. We encompass all thought chains in the refined thought structure $\mathcal{S}^{\prime}$. First, we use gradient information to recalculate and select key nodes of each chain. For each key node $\mathcal{T}_{i_{key}}$, we calculate a weight $w_{i_{key}}$ based on its impact on the solution as:
@@
where $L_{i_{key}}$ represents the self-information loss at node $\mathcal{T}_{i_{key}}$, which reflects the model’s confidence and potential errors in its prediction at that node. The contribution of each key node, denoted as $\mathbf{v_{i_{key}}}$, is the value vector that quantifies the influence of the node on the overall reasoning process. This vector can include various factors such as the relevance of the node’s content to the task or the accuracy of its inference. Then, for the reasoning task $\mathcal{Q}$, we compute the collaborative reasoning score by summing the contributions of all key nodes across all chains as:
$C(\mathcal{Q})=\sum_{i=1}^{N}\sum_{j=1}^{K_{i}}w_{i_{key}}\cdot\mathbf{v}_{ij}$ (14)
This process integrates the weighted contributions for a reasoning score that involves the individual node’s direct impact and its significance within the context of the entire chain. The final decision $D$ for the task $\mathcal{Q}$ is selected based on the highest collaborative reasoning score via
@@
LM-as-a-Judge
This strategy is suitable for reasoning tasks that require detailed interpretation and judgment, applicable to both gradient-available models and black-box models. Here, the model $\mathcal{L}$ acts as a judge to score each thought chain’s output based on its assessment of reasoning coherence, prediction confidence, or relevance to the task. Based on the evaluation, the model employs a voting mechanism to determine the final output:
$\text{Score}_{i}=\text{Evaluate}_{\mathcal{L}}(C_{q}),\quad D=\arg\max_{q\in\mathcal{Q}}C(q)$ (16)
@@
Experiments
Task Game of 24-GPT Creative Writing-GPT
Success Rate (%) Soundness Innovation Coherence Expression Overall
CoT 13.30 5.26 5.24 5.03 4.95 5.12
CoT-SC 46.70 5.41 5.17 5.17 5.16 5.23
ToT 52.70 5.90 5.36 5.40 5.43 5.52
RATT 41.30 6.02 6.22 5.69 5.67 5.90
TSE 74.00 6.19 6.55 5.74 5.64 6.03
Task Game of 24-Llama Creative Writing-Llama
Success Rate (%) Soundness Innovation Coherence Expression Overall
CoT 1.33 4.63 3.98 4.25 4.20 4.27
CoT-SC 5.33 4.84 4.35 4.51 4.46 4.54
ToT 8.00 4.86 4.95 4.78 4.65 4.81
RATT 8.67 5.10 4.97 4.89 4.97 4.98
TSE 18.75 5.03 5.11 5.05 5.15 5.09
@@
Table 1: Overall performance on reasoning tasks based on GPT-4o-mini and Llama. Bold and underline indicate the best and the second-best results.
Experimental Settings
Implementation Details and Baselines. We conduct experiments on three reasoning tasks that require mathematical, hierarchical, and comprehensive reasoning abilities to evaluate the effectiveness of our TSE method. We conduct these experiments utilizing the GPT-4o-mini (OpenAI 2024) and Llama-3.1-8B-Instruct (Hugging Face 2024). Unless otherwise specified, we generate five parallel thought chains with a depth of 5 for each question and use this structure as the basis for new node generation. The temperature for all models is set to the default value of 0.7, with a maximum token limit of 50. All tasks are performed on an NVIDIA 4090 GPU. We apply our method to the simplest multi-chain thought structure and compare the results with several baseline methods, including CoT, Vanilla CoT-SC, ToT, and RATT. This comparison aims to illustrate how our method enhances the thought structure compared to existing approaches.
@@
Task Description. We evaluate TSE and the baseline methods on three task-specific reasoning datasets: (1) Game of 24 (Yao et al. 2023), a mathematical challenge to use the four basic arithmetic operations to make four given numbers equal 24. The task requires the language models to combine multiple operations to achieve a target outcome, which can evaluate arithmetic reasoning and logical problem-solving capabilities. (2) Mini Crosswords (Yao et al. 2023), a game of 5×5 mini crosswords and each input includes the 5 horizontal and 5 vertical clues. To solve this task, the model requires deeper exploration and strategic integration of linguistic clues, allowing us to understand how effectively the evaluated model can expand traditional solution paths and uncover new insights within a complex search space; (3) Creative Writing (Yao et al. 2023), a task to construct a coherent passage with four paragraphs, each ending with one of four given sentences.This task compels LLMs to generate imaginative text with logically sound and contextually rich. Detailed information is provided in Appendix.
Overall Performance
@@
The overall performance is reported in Table 1 and Figure 3. We can see our TSE consistently outperforms compared baseline methods across different tasks on most of the metrics. The effectiveness of TSE is attributed to expanding thought structures and exploring solution space to contribute to different aspects of reasoning. Then, we analyze each task separately as follows.
@@
Game of 24. As shown in Table 1, the TSE-refined CoT-SC method significantly outperforms other thought structures, including vanilla CoT-SC, with an improvement of $58.56\%$, ToT by $40.50\%$, and RATT by $79.04\%$. Additionally, our enhancement of the basic CoT-SC structure achieves accuracy on GPT-4o-mini that matches the performance of the more complex ToT ($b=5$) implemented on GPT-4 (Yao et al. 2024). These results highlight the substantial improvements in reasoning capabilities brought about by our method.
@@
Mini Crossword. In this task, we evaluate performance based on two metrics: the proportion of correct letters (out of 25 per game) and the proportion of successfully solved games. As shown in Figure 3(b), the CoT-SC with TSE achieves an impressive accuracy rate of $82.4\%$, significantly outperforming vanilla CoT-SC (by $30.2\%$), surpassing ToT by $6.9\%$. However, for the Game metric, RATT performs slightly better than our method, as it can leverage external knowledge through RAG. Specifically, vanilla CoT-SC suffers from limited exploration of thought paths, which reduces its ability to consistently generate accurate answers. These results show that TSE enhances the model’s depth and diversity in exploring potential solutions and generates more accurate and coherent outputs.
@@
Creative Writing. As shown in Table 1, the evaluation criteria included Soundness, Innovation, Content Coherence of Reasoning, and Clarity of Expression, applying consistently across all evaluations. We rate each metric on a scale from 0 to 10 in 0.5-point increments, and calculate the overall performance score as the average of these four dimensions. A detailed figure can be found in the Appendix.
Ablation Study
Key Node Selection.
@@
To validate the importance of key node selection, we compare (1) the original CoT-SC, (2) the random selection of one node per chain for generation, and (3) our key node selection method. As shown in Table 3, our method significantly improves on Game of 24 compared to original and random approaches. In Creative Writing, particularly for innovation, our method also demonstrates notable enhancements, proving its effectiveness in selecting nodes for generation. The superior performance of our key node selection method is primarily due to its ability to identify nodes containing crucial information for problem-solving. The model only utilizes the most influential information and thus avoids redundancy, thus enhancing the model’s efficiency and creativity.
Connection Node Selection.
We then test the necessity of connection node selection by comparing (1) randomly selecting key nodes to connect, (2) selecting the key nodes based on lower inference layers, and (3) our method. As shown in Table 3, our method is significantly superior to random selection and layer-based selection on overall reasoning performance, particularly in coherence. This is because random selection lacks specificity and cannot guarantee that the chosen nodes will effectively support the content of new nodes. Although layer-based selection considers the structural hierarchy of information, it does not necessarily reflect the actual importance or applicability of the information. Our method analyzes relationships and reasoning paths between nodes, thus maintaining and strengthening the information flow’s coherence and depth.
Collaborative Reasoning.
In this experiment, we test the importance of the collaboration method in TSE. We compare four collaboration methods: (1) aggregating answers through majority voting, (2) randomly sampling partial thought chains, (3) using only the outcomes from newly generated chains, and (4) implementing the TSE collaborative method. As shown in Table 3, TSE consistently outperforms other methods in both tasks. Meanwhile, using only new chains also achieves a high success rate in Game of 24, demonstrating their potential to offer innovative solutions. However, since these new chains lack integration with original chain information, their performance is still weaker than that of TSE’s comprehensive collaboration methods. In Creative Writing, using only new chains proved better than majority voting. This suggests that while new chains can offer innovative content, they may lack coherence and depth without sufficient integration with the original chains.
Method GPT Llama
Success Rate (%) Overall Coherence Success Rate (%) Overall Coherence
@@
Random 62.8 5.91 5.66 9.52 4.82 4.85
Layer-based 67.3 5.95 5.69 6.12 4.84 4.88
TSE 74.0 6.03 5.74 18.75 5.09 5.05
@@
Majority-Vote 46.7 5.81 5.87 10.0 4.85 4.85
Random Sampling 27.3 5.90 5.89 12.5 4.92 4.90
New Chains Only 64.0 5.95 5.87 18.06 5.11 5.06
TSE 74.0 6.03 5.90 18.75 5.09 5.05
@@
Strategy Analysis
Impact of Key Node Selection Methods.
In this experiment, we compare the impact of two key node selection methods—gradient-based and self-prompting—on Llama’s CoT-SC reasoning performance. As shown in Figure 4, both methods achieve comparable results in the Game of 24, while self-prompting performs better in Creative Writing. This indicates both approaches effectively select critical nodes, but self-prompting additionally helps the model revisit overlooked thought connections and blind spots. Consequently, the model generates more reliable information and produces more creative content. Thus, for tasks demanding higher creativity, self-prompting can encourage the model to move beyond conventional thought patterns and foster innovative solutions.
@@
(a) Performance by letter.
(b) Performance by game.
Figure 3: Performance of GPT-4o-mini on Mini CrossWord.
Figure 4: Performance of key node selection methods.
Comparison of Connection Node Selection Methods.
We then compare gradient comparison and semantic relevance for selecting connection nodes when generating new reasoning paths on Llama. As shown in Table 3, semantic relevance selection achieves better performance on both Game of 24 and coherence in Creative Writing. The difference possibly arises from two main reasons: (1) Nodes with larger relative gradients do not necessarily have a stronger influence on the content when generating new nodes. (2) Compared to gradient-based selection, the semantic relevance method more effectively captures the actual content and meaning relationships between nodes, leading to more suitable connection points for new nodes. This demonstrates that considering semantic relevance when selecting connection points provides more effective support for reasoning performance.
Comparison of Collaborative Reasoning Methods.
In this experiment, we compare the performance of CWS and LM-as-a-Judge. As shown in Table 3, in task Game of 24, which requires dense logic and has well-defined objectives, the LM-as-a-Judge method accurately and quickly identifies the correct answers by directly evaluating each thought chain and selecting the optimal solution. This method relies on the model’s ability to assess individual outputs, allowing for rapid selection of the best solution—especially useful in tasks that require precise calculations and quick responses. In task Creative Writing, which requires innovation and diversity, CWS evaluates and weights the contributions of each thought chain. The final output represents a blend of multiple perspectives and enhances the richness and depth of the narrative significantly. We further analyze the efficacy and provide a case study to illustrate the effectiveness of TSE generation and collaboration in the Appendix.
Conclusion
@@ Appendix A Experimental Settings
Task Description
We evaluate TSE and the baseline methods on three reasoning datasets with specific tasks: (1) Game of 24, a mathematical challenge whose objective is to use the four basic arithmetic operations to make four given numbers equal 24. The task requires the language models to combine multiple operations to achieve a target outcome, which can evaluate arithmetic reasoning and logical problem-solving capabilities. (2) Mini Crosswords, a game of 5×5 mini crosswords and each input includes the 5 horizontal and 5 vertical clues. The expected output is a completed $5×5$ crossword board containing 25 letters. To solve this task, the model requires deeper exploration and strategic integration of linguistic clues, allowing us to understand how effectively the evaluated model can expand traditional solution paths and uncover new insights within a complex search space; (3) Creative Writing, a task to construct a coherent passage with four paragraphs, each ending with one of four given sentences.This task compels LLMs to generate imaginative text withlogically sound, and contextually rich. To ensure a rigorous and comprehensive assessment, we conduct half of the evaluations using GPT-4 (OpenAI 2023), while the other half are performed by a panel of expert human annotators. In this task, we select 100 open-ended questions from several prompts listed on Reedsy. com [^2] https://blog.reedsy.com/creative-writing-prompts/“#:“˜:text=When“%20the“%20idea“%20to“%20start“%20a“%20weekly“%20newsletter as our input.
Thought Chain Efficacy
@@
Creative Writing
In Figure 5 we provide detailed results on task Creative Writing. TSE method outperforms other methods in three aspects and is only slightly weaker than RATT with RAG.